OpenStudy (sparklestaraa):

Jess plans to increase the amount of money she saves each month. She can increase her savings in the following ways:

OpenStudy (sparklestaraa):

@iLorita

OpenStudy (sparklestaraa):

@campbell_st

OpenStudy (sparklestaraa):

@iLorita SEEEEEEEE

OpenStudy (anonymous):

@Albany_Goon

OpenStudy (anonymous):

@Phebe

OpenStudy (anonymous):

heyy wassup geek squaddddd lml

OpenStudy (phebe):

im hea

OpenStudy (sparklestaraa):

wassupppp :D

OpenStudy (phebe):

wussup

OpenStudy (sparklestaraa):

not much enduring the math -,-

OpenStudy (phebe):

lol i rly dont like math it sucks too me ~just my opinion~

OpenStudy (anonymous):

ok srry bout that lemme c here

OpenStudy (sparklestaraa):

same here

OpenStudy (sparklestaraa):

anyone have an answer xd

OpenStudy (anonymous):

Basically, what you should do is, begin with option 1 and divide each set of options consecutively starting with the 5th and 4th options: 29.28/26.62 = 1.1 Then the 4th and 3rd options 26.62/24.2 = 1.1 Then the 3rd and 2nd options: 24.2/22 = 1.1 Then the 2nd and 1st options: 22/20 = 1.1 Notice that the rate of increase is the same for each So we know that at least option 1 is correct.